# PIF-Net: Ill-Posed Prior Guided Multispectral and Hyperspectral Image Fusion via Invertible Mamba and Fusion-Aware LoRA

### v1

@@ Introduction
Multispectral and Hyperspectral Image Fusion (MHIF) aims to effectively combine the rich spatial details from multispectral images (MSI) with the continuous and abundant spectral information from hyperspectral images (HSI), generating high-quality images that possess both high spatial resolution and high spectral resolution. By integrating data from these two modalities, MHIF seeks to overcome the limitations of single sensors in either spatial or spectral resolution, enhancing spectral fidelity and spatial detail representation in the fused images. This, in turn, better supports vision-based downstream applications such as image classification (Chen et al. 2025; Ding et al. 2025; Cao et al. 2024), object detection (Qin et al. 2024; Liu et al. 2024a; He et al. 2023), and other practical tasks (Sethy et al. 2022; Bi et al. 2024; Bhargava et al. 2024).
@@
Most fusion methods based on single-branch (as Figure 1(a)) (Hu et al. 2022b; Ma et al. 2023; Deng et al. 2023) or dual-branch (as Figure 1(b)) (Peng et al. 2023; Hu et al. 2022c) architectures, despite showing improvements over traditional model-driven approaches (He et al. 2016; Vivone et al. 2014; Aiazzi, Baronti, and Selva 2007), still suffer from two fundamental limitations: (1) Significant differences in spectral characteristics, spatial resolution, and data distribution between MSIs and HSIs result in severe information misalignment and a substantial modality gap between the two sources. This constitutes a highly ill-posed problem that renders unified and effective feature modeling particularly challenging, thereby constraining further advances in fusion performance. (2) Existing fusion methods commonly fail to ensure reversible information transfer, leading to inevitable information loss during the fusion process. Such irreversible loss directly impairs spectral fidelity and the high-quality recovery of spatial details.
@@
To address the ill-posedness inherent in image restoration or image fusion, researchers have conducted extensive studies and explorations. For example, the Structure-Aware Deep Network (Fan et al. 2025; Lu, You, and Barnes 2018) enhances the model’s ability to capture geometric structures in images by introducing a structure-aware module, thereby significantly improving the visual quality of super-resolved images. The adoption of generative adversarial training (Creswell et al. 2018; Ma et al. 2020; Wang et al. 2018) further facilitates the generation of realistic high-resolution images with complex structures. Denoising diffusion models (Sun, Chen, and Messinger 2014; Saharia et al. 2022) have demonstrated excellent performance in detail restoration and noise suppression, producing more natural and realistic images while effectively mitigating the over-smoothing caused by the ill-posed nature of the problem. However, these methods typically rely on general structural priors and mainly focus on low-level fixed patterns, making it difficult to capture the semantic shifts introduced during fusion.
@@
Furthermore, the introduction of invertible neural networks (INNs) (Dinh, Krueger, and Bengio 2014; Xiao et al. 2020) offers a novel perspective for image restoration tasks. Unlike traditional unidirectional mapping, invertible networks learn bidirectionally consistent mappings, establishing a reversible information flow between image degradation and restoration (Kingma and Dhariwal 2018; Lugmayr et al. 2020), thereby theoretically preventing information loss. This mechanism not only helps alleviate the ill-posedness in the restoration process but also significantly enhances the model’s capability to transfer information across different modalities, making it particularly suitable for the complex cross-modal modeling demands in multi-source image fusion. Nevertheless, effectively leveraging the invertible structure to address the issue of information misalignment caused by modality inconsistencies remains a significant challenge.
@@
To address the above issues, this paper proposes an advanced fusion framework guided by ill-posed priors, as shown in Figure 1(c), called PIF-Net. This framework ingeniously integrates the invertible visual state-space model Mamba with a lightweight Fusion-Aware Multi-Head Low-Rank Adaptation (FAM-LoRA) module , achieving high-fidelity and efficient multimodal image fusion. PIF-Net consists of two complementary branches: the frequency domain branch precisely separates high- and low-frequency components using a two-dimensional Haar wavelet transform, combined with the invertible Mamba module to enhance bidirectional information flow, ensuring lossless extraction of global spatial features and structural fidelity; the spatial domain branch leverages low-frequency global guidance and the lightweight LoRA module to finely extract rich spatial texture information, incorporating Large-Kernel Attention (LKA) (Wang et al. 2023; Liu et al. 2022; Peng et al. 2017) and Squeeze-and-Excitation (SE) (Hu, Shen, and Sun 2018) mechanisms to significantly improve cross-modal semantic modeling. Furthermore, PIF-Net introduces guided feature consistency loss and volumetric-aware loss to strengthen the consistency between the dual-modal features, ensuring high precision in modality alignment and feature fusion with semantic coherence. Extensive experimental results demonstrate that PIF-Net outperforms current state-of-the-art methods across multiple hyperspectral datasets, achieving outstanding fusion quality and computational efficiency.
@@
In the field of multispectral and hyperspectral image fusion (MHIF), traditional methods such as pansharpening (Wu et al. 2021; Aiazzi, Baronti, and Selva 2007) and multiscale decomposition (Yang et al. 2017) offer good interpretability but heavily rely on prior knowledge, making them prone to failure in complex scenarios. Bayesian approaches (Yang et al. 2018; Aiazzi et al. 2006) can effectively model the fusion process but require careful parameter tuning. In recent years, deep learning methods have significantly advanced MHIF. Early convolutional neural network (CNN)-based models, such as MHFnet (Xie et al. 2022) and HSRnet (Hu et al. 2022c), perform well under general conditions but struggle to capture long-range dependencies, leading to degraded performance in complex cases. To address this, Transformer-based Fusformer (Hu et al. 2022b) was introduced, which greatly improves fusion quality but at a higher computational cost. Other approaches like PSRT (Deng et al. 2023), U2Net (Peng et al. 2023), and 3DT-Net (Ma et al. 2023) bring innovations in architecture design but still face limitations in fusion capability, high training costs, or increased model complexity. SMGU-Net (Yan et al. 2025) attempts to combine traditional methods with deep learning to enhance cross-modal interactions, yet there remains substantial room for improvement in modeling complex nonlinear cross-modal relationships.
@@
Invertible Neural Networks (INNs) are a key component of normalizing flow models. Since NICE (Dinh, Krueger, and Bengio 2014) introduced additive coupling layers, methods like RealNVP (Dinh, Sohl-Dickstein, and Bengio 2017) and Glow (Kingma and Dhariwal 2018) have improved INNs, enabling efficient high-resolution image generation. Beyond generative tasks, INNs can losslessly preserve features and save memory in classification. They also show broad potential in image colorization (Ardizzone et al. 2019), steganography (Zhang, Cuesta-Infante, and Veeramachaneni 2019), image rescaling, and super-resolution (Xiao et al. 2020; Lugmayr et al. 2020). However, INNs remain underexplored in image fusion. For example, CDDFuse (Zhao et al. 2023) uses convolution-based invertible modules for detail extraction but fails to fully exploit spatial guidance from auxiliary modalities to improve fusion quality.
@@
The Residue Channel Prior (RCP) (Yi et al. 2021; Liao et al. 2025), as an effective ill-posed prior in image restoration, exhibits spatial invariance while preserving contextual information. Inspired by this, we propose a dynamic Ill-Posed Residual Prior Extraction Module (IPRPEM) tailored for the MHIF task. This module is designed to adaptively capture spectral-domain ill-posed characteristics while preserving spatially shifted contextual features. Specifically, it extracts global ill-posed prior information from shallow feature differences in an adaptive manner, enhancing the discriminative capacity of subsequent feature fusion. The computation is formulated as:
{$C_{R}&=\max_{c\in D}\bar{\mathbf{X}}_{s}-\min_{c\in D}\bar{\mathbf{X}}_{s},\\ \mathbf{X}_{R}&=\text{ReLU}\big{(}\text{Conv}(\mathbf{X}_{s}-\mathbf{Y}_{s})\big{)}\cdot C_{R},\\ \mathbf{X}_{R}&=\text{ReLU}\big{(}\text{Conv}(\mathbf{X}_{R})\big{)},\end{aligned}\right.$ (2)
@@
Figure 3: Overview of the Invertible Mamba Block. The block takes low-frequency components $\mathbf{X}^{L}_{i}$ and high-frequency components $\mathbf{X}^{H}_{i}$ as input, and uses an affine coupling mechanism based on lightweight Segmented Spectral Mamba Modules (SSMM) to enable efficient bidirectional interaction and fusion of spectral features. The operations Split and Concat refer to feature division and aggregation along the channel dimension. SS2D refers to the 2D State Space Module in VMamba (Liu et al. 2024c).
@@
CNNs (Hu et al. 2022c; Peng et al. 2023) struggle to capture long-range dependencies. Transformers (Hu et al. 2022b) offer strong global modeling but suffer from quadratic complexity as feature size grows, reducing efficiency. To balance performance and cost, we adopt Mamba (Liu et al. 2024c; Gu and Dao 2023), a state-space model with linear complexity. Mamba captures global context efficiently and expressively. However, most Mamba-based models (Guo et al. 2024, 2025; Li, Wang, and Xu 2025) still rely on irreversible designs, which cause feature ambiguity and information loss in image restoration. To solve this, we propose the Invertible Mamba Block, which enables bidirectional interaction between low- and high-frequency spectral features (see Figure 3(a)).
@@
$\mathbf{X}_{L}^{i+1}$ $=\mathbf{X}_{L}^{i}\circ\exp\left[I_{2}(\mathbf{X}_{H}^{i})\right]+I_{3}(\mathbf{X}_{H}^{i}).$
Here, $\circ$ denotes element-wise multiplication. $I_{1}$, $I_{2}$, and $I_{3}$ are SSMM-based subnetworks designed to model nonlinear interactions between different spectral components. Thanks to the invertible architecture, the module enables efficient information fusion during the forward pass.
Notably, to ensure the flexibility of invertible blocks in fusion tasks, we employ a gradient-guided invertibility strategy that heuristically facilitates backward interaction of spectral features. Instead of using traditional rescaling schemes (Xiao et al. 2020), we design a tailored invertible loss function (see Eq. (4)).
@@
Brovey $\times$2 29.2704 0.8751 8.1522 11.0668 32.5312 0.8431 8.4829 5.7936 27.8390 0.7958 6.2207 6.0319
GSA $\times$2 37.4039 0.9627 4.1328 7.0972 33.2998 0.8630 9.3221 5.4163 28.7031 0.7945 6.6719 6.4978
HSRnet $\times$2 53.0324 0.9964 1.9876 1.4615 40.8710 0.9805 3.4652 1.9588 40.9947 0.9685 3.6476 2.2422
Fusformer $\times$2 53.8104 0.9967 1.9414 1.4521 42.1020 0.9888 3.1807 1.9238 41.8442 0.9727 3.2919 2.0589
PSRT $\times$2 53.8821 0.9970 1.8994 1.3537 42.3937 0.9901 2.8675 1.8847 41.9956 0.9790 2.2518 1.8971
U2Net $\times$2 53.9657 0.9974 1.8824 1.2972 42.7243 0.9920 2.7825 1.8308 42.2347 0.9813 2.1596 1.4162
3DT-Net $\times$2 53.9155 0.9988 1.8222 1.3098 43.1856 0.9947 2.5811 1.7677 41.2224 0.9807 2.3080 1.6433
SMGU-Net $\times$2 52.7751 0.9985 1.8869 1.2842 43.1647 0.9946 2.6137 1.7902 39.3881 0.9775 2.8193 1.8930
@@
Brovey $\times$4 27.5923 0.8325 6.2724 15.1083 30.1662 0.7745 11.5137 7.9550 25.5351 0.7143 8.8247 7.9284
GSA $\times$4 31.2757 0.8985 6.0018 10.3126 30.9217 0.7605 12.3469 8.5070 28.5528 0.6295 9.5707 10.0441
HSRnet $\times$4 49.3548 0.9901 2.4509 1.9242 40.4509 0.9835 3.8988 2.9216 37.2608 0.9540 3.7748 3.1791
Fusformer $\times$4 50.1466 0.9909 2.3864 1.8927 41.1338 0.9895 3.4892 2.3436 37.7936 0.9558 3.8403 2.9017
PSRT $\times$4 50.5377 0.9919 2.2097 1.7832 41.5616 0.9882 3.2434 2.1695 38.2794 0.9645 3.3579 2.4509
U2Net $\times$4 50.5061 0.9929 2.1852 1.7317 41.7714 0.9889 3.1039 1.9978 38.7681 0.9700 3.0176 2.2220
3DT-Net $\times$4 48.1940 0.9961 2,4262 2.1546 42.6873 0.9931 3.0318 1.9952 38.1864 0.9744 2.8405 2.3481
SMGU-Net $\times$4 51.3382 0.9981 2.1080 1.5762 42.9130 0.9937 2.8873 1.8979 38.9201 0.9761 2.9800 1.9973
@@
Brovey $\times$8 26.0978 0.7989 8.1166 18.1408 27.0096 0.7217 16.0689 11.4050 22.7245 0.6497 13.4666 10.7571
GSA $\times$8 28.1606 0.8609 8.0270 13.4385 26.2180 0.6829 17.6147 11.9947 21.2674 0.5245 14.1969 14.3429
HSRnet $\times$8 47.7262 0.9902 2.6291 2.3439 30.6384 0.9550 6.0210 5.9721 30.1921 0.8956 6.0662 4.9901
Fusformer $\times$8 47.8910 0.9919 2.5806 2.2506 31.8052 0.9589 5.7651 5.4997 30.4891 0.9011 5.9990 4.9715
PSRT $\times$8 48.6457 0.9942 2.5291 2.0439 32.6384 0.9650 5.1151 5.1257 31.7671 0.9189 5.6891 4.8965
U2Net $\times$8 48.8911 0.9959 2.4006 1.9506 32.8052 0.9676 4.9910 4.9711 31.9211 0.9210 5.5451 4.6712
3DT-Net $\times$8 44.9324 0.9936 2,9072 2.6482 34.1538 0.9688 5.0265 4.8583 29.2755 0.9302 5.8975 5.0295
SMGU-Net $\times$8 49.6582 0.9975 2.3260 1.7872 34.2202 0.9699 4.6268 4.8072 29.7848 0.9417 5.4906 4.4843
@@
Best Value - +$\infty$+ ∞ 1 0 0 +$\infty$+ ∞ 1 0 0 +$\infty$+ ∞ 1 0 0
Figure 4: Overview of the FAM-LoRA module that fuses the main spatial feature $\mathbf{Y}_{i}$ and auxiliary guidance $\mathbf{X}^{R}_{i}$ through channel transformation, LKA, SE attention, and multi-head LoRA, achieving efficient and accurate semantic fusion. The first part of LoRA’s parameters is initialized with a standard Gaussian distribution, while the second part is initialized to zero. The operations Split and Concat refer to feature division and aggregation along the channel dimension.
@@
In recent years, significant breakthroughs have been made in fine-tuning large language models (LLMs) based on LoRA (Hu et al. 2022a; Liu et al. 2024b), followed by rapid developments of LoRA models in the vision domain (Zhong et al. 2024; Talla et al. 2017). LoRA efficiently approximates large-scale parameter models through the product of two low-rank matrices, significantly reducing the number of trainable parameters and providing an efficient and concise representation method for feature fusion. Inspired by this, we designed a Fusion-Aware Multi-head Low-Rank Adaptation module (named FAM-LoRA) as shown in Figure 4, which achieves efficient semantic interaction during feature fusion while maintaining very low parameter count and computational complexity.
Specifically, FAM-LoRA takes the main spatial feature $\mathbf{Y}_{i}$ and the auxiliary guidance feature $\mathbf{X}^{R}_{i}$ as inputs. The main feature $\mathbf{Y}_{i}$ first passes through four consecutive $\times 1$ 1 convolutional layers for channel compression and transformation. Then, the module integrates LKA and SE mechanisms, significantly enhancing spatial contextual awareness and channel-wise selectivity. After the second attention operation, the gradients of the LKA and SE modules are frozen, and the auxiliary feature $\mathbf{X}^{R}_{i}$ is injected into the input of the SE module, further guiding the attention mechanism to focus on critical semantic information, thereby improving the precision and expressiveness of feature fusion. To enhance multi-scale feature representation, we adopt a multi-head LoRA design that evenly splits the channel dimension into four parts, applying low-rank adaptation to each separately.
@@
$\mathcal{L}=\underbrace{\|\hat{\mathbf{Z}}-\mathbf{Z}\|_{1}}_{\mathcal{L}_{1}}+\lambda_{\mathrm{inv}}\cdot\underbrace{\log\left|\det\left(\frac{\partial\bar{\mathbf{Z}}}{\partial\mathbf{X}}\right)\right|}_{\mathcal{L}_{inv}}+\lambda_{\mathrm{cos}}\cdot\underbrace{\left(1-\cos(\bar{\mathbf{Z}},\hat{\mathbf{Z}})\right)}_{\mathcal{L}_{cos}}$ (4)
@@
Second, the log-determinant of the Jacobian of the affine transformation, $\mathcal{L}_{inv}=\log\left|\det\left(\frac{\partial\bar{\mathbf{Z}}}{\partial\mathbf{X}}\right)\right|$, serves as an invertibility regularization term to ensure information preservation and numerical stability during the mapping process (Kingma and Dhariwal 2018; Lugmayr et al. 2020), encouraging the model to learn stable and invertible feature transformations. Finally, the cosine similarity loss $\mathcal{L}_{cos}=1-\cos(\bar{\mathbf{Z}},\hat{\mathbf{Z}})$ enhances angular consistency between different feature representations, improving semantic feature fusion.
@@
We conduct experiments on three widely-used hyperspectral image datasets: Chikusei (Yokoya and Iwasaki 2016), Pavia University (PaviaU), and Houston. The Chikusei dataset contains 2517×23352517$\times$ 23352517 × 2335 pixels with 128 spectral bands. We select a 1000×20001000$\times$ 20001000 × 2000 region from the top-left corner for training, while the remaining area is divided into non-overlapping $\times 680$ 680 patches for testing. The PaviaU dataset includes 103 spectral bands and has a spatial size of 610×340610$\times$ 340610 × 340 pixels. The top $\times 340$ 340 region is reserved for testing, with the rest used for training. The Houston dataset contains 349×1905349$\times$ 1905349 × 1905 pixels across 144 spectral bands; the left $\times 349$ 349 area is used for testing, and the rest for training. Following the protocol in (Wei et al. 2015), HRMSI are simulated, and LRHSI are generated by applying Gaussian blur (kernel size $\times 3$ 3, standard deviation 0.5) followed by downsampling. The training data are cropped into $\times 64$ 64 patches as ground truth, paired with corresponding LR-HSI patches ($\times 16$ 16) and HRMSI patches as inputs.
@@
For quantitative evaluation, we adopt four commonly-used metrics: Peak Signal-to-Noise Ratio (PSNR), Spectral Angle Mapper (SAM) (Yuhas, Goetz, and Boardman 1992), Relative Global Dimensional Synthesis Error (ERGAS) (Wald 2002), and Structural Similarity Index (SSIM) (Wang et al. 2004).

### v2

@@ Introduction
Multispectral and Hyperspectral Image Fusion (MHIF) aims to effectively combine the rich spatial details from multispectral images (MSI) with the continuous and abundant spectral information from hyperspectral images (HSI), generating high-quality images that possess both high spatial resolution and high spectral resolution. By integrating data from these two modalities, MHIF seeks to overcome the limitations of single sensors in either spatial or spectral resolution, enhancing spectral fidelity and spatial detail representation in the fused images. This, in turn, better supports vision-based downstream applications such as image classification (class_01; class_02; class_03), object detection (od_02; od_01; od_03), and other practical tasks (agriculture; diagnosis; application).
@@
Most fusion methods based on single-branch (as Figure 1(a)) (Fusformer; 3DT-Net; PSRT) or dual-branch (as Figure 1(b)) (U2Net; HSRnet) architectures, despite showing improvements over traditional model-driven approaches (low_rank01; Brovey; GSA), still suffer from two fundamental limitations: (1) Significant differences in spectral characteristics, spatial resolution, and data distribution between MSIs and HSIs result in severe information misalignment and a substantial modality gap between the two sources. This constitutes a highly ill-posed problem that renders unified and effective feature modeling particularly challenging, thereby constraining further advances in fusion performance. (2) Existing fusion methods commonly fail to ensure reversible information transfer, leading to inevitable information loss during the fusion process. Such irreversible loss directly impairs spectral fidelity and the high-quality recovery of spatial details.
@@
To address the ill-posedness inherent in image restoration or image fusion, researchers have conducted extensive studies and explorations. For example, the Structure-Aware Deep Network (structure_02; structure_01) enhances the model’s ability to capture geometric structures in images by introducing a structure-aware module, thereby significantly improving the visual quality of super-resolved images. The adoption of generative adversarial training (gan; pan-gan; esrgan) further facilitates the generation of realistic high-resolution images with complex structures. Denoising diffusion models (pan-diffusion; sr3) have demonstrated excellent performance in detail restoration and noise suppression, producing more natural and realistic images while effectively mitigating the over-smoothing caused by the ill-posed nature of the problem. However, these methods typically rely on general structural priors and mainly focus on low-level fixed patterns, making it difficult to capture the semantic shifts introduced during fusion.
@@
Furthermore, the introduction of invertible neural networks (INNs) (2014NICE; xiao2020invertible) offers a novel perspective for image restoration tasks. Unlike traditional unidirectional mapping, invertible networks learn bidirectionally consistent mappings, establishing a reversible information flow between image degradation and restoration (Gflow; lugmayr2020srflow), thereby theoretically preventing information loss. This mechanism not only helps alleviate the ill-posedness in the restoration process but also significantly enhances the model’s capability to transfer information across different modalities, making it particularly suitable for the complex cross-modal modeling demands in multi-source image fusion. Nevertheless, effectively leveraging the invertible structure to address the issue of information misalignment caused by modality inconsistencies remains a significant challenge.
@@
To address the above issues, this paper proposes an advanced fusion framework guided by ill-posed priors, as shown in Figure 1(c), called PIF-Net. This framework ingeniously integrates the invertible visual state-space model Mamba with a lightweight Fusion-Aware Multi-Head Low-Rank Adaptation (FAM-LoRA) module , achieving high-fidelity and efficient multimodal image fusion. PIF-Net consists of two complementary branches: the frequency domain branch precisely separates high- and low-frequency components using a two-dimensional HWT, combined with the invertible Mamba module to enhance bidirectional information flow, ensuring lossless extraction of global spatial features and structural fidelity; the spatial domain branch leverages low-frequency global guidance and the lightweight LoRA module to finely extract rich spatial texture information, incorporating Large-Kernel Attention (LKA) (lka1; lka2; lka3) and Squeeze-and-Excitation (SE) (SENet) mechanisms to significantly improve cross-modal semantic modeling. Furthermore, PIF-Net introduces guided feature consistency loss and volumetric-aware loss to strengthen the consistency between the dual-modal features, ensuring high precision in modality alignment and feature fusion with semantic coherence. Extensive experimental results demonstrate that PIF-Net outperforms current state-of-the-art methods across multiple hyperspectral datasets, achieving outstanding fusion quality and computational efficiency.
@@
In the field of multispectral and hyperspectral image fusion (MHIF), traditional methods such as pansharpening (pan15; GSA) and multiscale decomposition (deco) offer good interpretability but heavily rely on prior knowledge, making them prone to failure in complex scenarios. Bayesian approaches (18; 19) can effectively model the fusion process but require careful parameter tuning. In recent years, deep learning methods have significantly advanced MHIF. Early convolutional neural network (CNN)-based models, such as MHFnet (MHFnet) and HSRnet (HSRnet), perform well under general conditions but struggle to capture long-range dependencies, leading to degraded performance in complex cases. To address this, Transformer-based Fusformer (Fusformer) was introduced, which greatly improves fusion quality but at a higher computational cost. Other approaches like PSRT (PSRT), U2Net (U2Net), and 3DT-Net (3DT-Net) bring innovations in architecture design but still face limitations in fusion capability, high training costs, or increased model complexity. SMGU-Net (SMGU-Net) attempts to combine traditional methods with deep learning to enhance cross-modal interactions, yet there remains substantial room for improvement in modeling complex nonlinear cross-modal relationships.
@@
Invertible Neural Networks (INNs) are a key component of normalizing flow models. Since NICE (2014NICE) introduced additive coupling layers, methods like RealNVP (RealNVP) and Glow (Gflow) have improved INNs, enabling efficient high-resolution image generation. Beyond generative tasks, INNs can losslessly preserve features and save memory in classification. They also show broad potential in image colorization (ardizzone2019guided), steganography (zhang2019steganogan), image rescaling, and super-resolution (xiao2020invertible; lugmayr2020srflow). However, INNs remain underexplored in image fusion. For example, CDDFuse (zhao2023cddfuse) uses convolution-based invertible modules for detail extraction but fails to fully exploit spatial guidance from auxiliary modalities to improve fusion quality.
@@
The Residue Channel Prior (RCP) (RCP_01; RCP_02), as an effective ill-posed prior in image restoration, exhibits spatial invariance while preserving contextual information. Inspired by this, we propose a dynamic Ill-Posed Residual Prior Extraction Module (IPRPEM) tailored for the MHIF task. This module is designed to adaptively capture spectral-domain ill-posed characteristics while preserving spatially shifted contextual features. Specifically, it extracts global ill-posed prior information from shallow feature differences in an adaptive manner, enhancing the discriminative capacity of subsequent feature fusion. The computation is formulated as:
$\left\{\begin{aligned} C_{R}&=\max_{c\in D}\bar{\mathbf{X}}_{s}-\min_{c\in D}\bar{\mathbf{X}}_{s},\\ \mathbf{X}_{R}&=\text{ReLU}\big(\text{Conv}(\mathbf{X}_{s}-\mathbf{Y}_{s})\big)\cdot C_{R},\\ \mathbf{X}_{R}&=\text{ReLU}\big(\text{Conv}(\mathbf{X}_{R})\big),\end{aligned}\right$. (2)
@@
Figure 3: Illustration of the Invertible Mamba Block. The block takes low-frequency features $\mathbf{X}^{L}_{i}$ and high-frequency features $\mathbf{X}^{H}_{i}$ as input, and employs an affine coupling mechanism built on lightweight Segmented Spectral Mamba Modules (SSMM) to enable efficient bidirectional interaction and fusion of spectral information. Split and Concat denote channel-wise feature division and aggregation, respectively. SS2D refers to the 2D State Space Module in VMamba (liu2024vmamba).
@@
CNNs (HSRnet; U2Net) struggle to capture long-range dependencies. Transformers (Fusformer) offer strong global modeling but suffer from quadratic complexity as feature size grows, reducing efficiency. To balance performance and cost, we adopt Mamba (liu2024vmamba; gu2023mamba), a state-space model with linear complexity. Mamba captures global context efficiently and expressively. However, most Mamba-based models (guo2024mambair; guo2025mambairv2; li2025hsrmamba) still rely on irreversible designs, which cause feature ambiguity and information loss in image restoration. To solve this, we propose the Invertible Mamba Block, which enables bidirectional interaction between low- and high-frequency spectral features (see Figure 3(a)).
@@
$\mathbf{X}_{L}^{i+1}$ $=\mathbf{X}_{L}^{i}\odot\exp\left[I_{2}(\mathbf{X}_{H}^{i})\right]+I_{3}(\mathbf{X}_{H}^{i}).$
Here, $\odot$ denotes element-wise multiplication. $I_{1}$, $I_{2}$, and $I_{3}$ are SSMM-based subnetworks designed to model nonlinear interactions between different spectral components. Thanks to the invertible architecture, the module enables efficient information fusion during the forward pass. Notably, to ensure the flexibility of invertible blocks in fusion tasks, we employ a gradient-guided invertibility strategy that heuristically facilitates backward interaction of spectral features. Instead of using traditional rescaling schemes (xiao2020invertible), we design a tailored invertible loss function (see Eq. (4)).
@@
Brovey (Brovey) $\times$2 29.2704 0.8751 8.1522 11.0668 32.5312 0.8431 8.4829 5.7936 27.8390 0.7958 6.2207 6.0319
GSA (GSA) $\times$2 37.4039 0.9627 4.1328 7.0972 33.2998 0.8630 9.3221 5.4163 28.7031 0.7945 6.6719 6.4978
HSRnet (HSRnet) $\times$2 53.0324 0.9964 1.9876 1.4615 40.8710 0.9805 3.4652 1.9588 40.9947 0.9685 3.6476 2.2422
Fusformer (Fusformer) $\times$2 53.8104 0.9967 1.9414 1.4521 42.1020 0.9888 3.1807 1.9238 41.8442 0.9727 3.2919 2.0589
PSRT (PSRT) $\times$2 53.8821 0.9970 1.8994 1.3537 42.3937 0.9901 2.8675 1.8847 41.9956 0.9790 2.2518 1.8971
U2Net (U2Net) $\times$2 53.9657 0.9974 1.8824 1.2972 42.7243 0.9920 2.7825 1.8308 42.2347 0.9813 2.1596 1.4162
3DT-Net (3DT-Net) $\times$2 53.9155 0.9988 1.8222 1.3098 43.1856 0.9947 2.5811 1.7677 41.2224 0.9807 2.3080 1.6433
SMGU-Net (SMGU-Net) $\times$2 52.7751 0.9985 1.8869 1.2842 43.1647 0.9946 2.6137 1.7902 39.3881 0.9775 2.8193 1.8930
@@
Brovey (Brovey) $\times$4 27.5923 0.8325 6.2724 15.1083 30.1662 0.7745 11.5137 7.9550 25.5351 0.7143 8.8247 7.9284
GSA (GSA) $\times$4 31.2757 0.8985 6.0018 10.3126 30.9217 0.7605 12.3469 8.5070 28.5528 0.6295 9.5707 10.0441
HSRnet (HSRnet) $\times$4 49.3548 0.9901 2.4509 1.9242 40.4509 0.9835 3.8988 2.9216 37.2608 0.9540 3.7748 3.1791
Fusformer (Fusformer) $\times$4 50.1466 0.9909 2.3864 1.8927 41.1338 0.9895 3.4892 2.3436 37.7936 0.9558 3.8403 2.9017
PSRT (PSRT) $\times$4 50.5377 0.9919 2.2097 1.7832 41.5616 0.9882 3.2434 2.1695 38.2794 0.9645 3.3579 2.4509
U2Net (U2Net) $\times$4 50.5061 0.9929 2.1852 1.7317 41.7714 0.9889 3.1039 1.9978 38.7681 0.9700 3.0176 2.2220
3DT-Net (3DT-Net) $\times$4 48.1940 0.9961 2.4262 2.1546 42.6873 0.9931 3.0318 1.9952 38.1864 0.9744 2.8405 2.3481
SMGU-Net (SMGU-Net) $\times$4 51.3382 0.9981 2.1080 1.5762 42.9130 0.9937 2.8873 1.8979 38.9201 0.9761 2.9800 1.9973
@@
Brovey (Brovey) $\times$8 26.0978 0.7989 8.1166 18.1408 27.0096 0.7217 16.0689 11.4050 22.7245 0.6497 13.4666 10.7571
GSA (GSA) $\times$8 28.1606 0.8609 8.0270 13.4385 26.2180 0.6829 17.6147 11.9947 21.2674 0.5245 14.1969 14.3429
HSRnet (HSRnet) $\times$8 47.7262 0.9902 2.6291 2.3439 30.6384 0.9550 6.0210 5.9721 30.1921 0.8956 6.0662 4.9901
Fusformer (Fusformer) $\times$8 47.8910 0.9919 2.5806 2.2506 31.8052 0.9589 5.7651 5.4997 30.4891 0.9011 5.9990 4.9715
PSRT (PSRT) $\times$8 48.6457 0.9942 2.5291 2.0439 32.6384 0.9650 5.1151 5.1257 31.7671 0.9189 5.6891 4.8965
U2Net (U2Net) $\times$8 48.8911 0.9959 2.4006 1.9506 32.8052 0.9676 4.9910 4.9711 31.9211 0.9210 5.5451 4.6712
3DT-Net (3DT-Net) $\times$8 44.9324 0.9936 2.9072 2.6482 34.1538 0.9688 5.0265 4.8583 29.2755 0.9302 5.8975 5.0295
SMGU-Net (SMGU-Net) $\times$8 49.6582 0.9975 2.3260 1.7872 34.2202 0.9699 4.6268 4.8072 29.7848 0.9417 5.4906 4.4843
@@
Best Value - $\infty$ 1 0 0 $\infty$ 1 0 0 $\infty$ 1 0 0
Figure 4: Illustration of the FAM-LoRA module that fuses the main spatial feature $\mathbf{Y}_{i}$ and auxiliary guidance $\mathbf{X}^{R}_{i}$ through channel transformation, LKA, SE attention, and multi-head LoRA, achieving efficient and accurate semantic fusion. The first part of LoRA’s parameters is initialized with a standard Gaussian distribution, while the second part is initialized to zero. The operations Split and Concat refer to feature division and aggregation along the channel dimension.
@@
Inspired by the success of LoRA in large language models and vision tasks (hu2022lora; dora; vis_lora_01; vis_lora_02), we propose a Fusion-Aware Multi-head Low-Rank Adaptation module (FAM-LoRA), as shown in Figure 4, which enables efficient semantic interaction during feature fusion with minimal parameters and computational overhead. Specifically, FAM-LoRA takes the main spatial feature $\mathbf{Y}_{i}$ and auxiliary guidance feature $\mathbf{X}^{R}_{i}$ as inputs. $\mathbf{Y}_{i}$ first passes through four consecutive $1\times 1$ convolutions for channel compression and transformation, followed by LKA and SE modules to enhance spatial context awareness and channel-wise selectivity. The gradients of LKA and SE are then frozen, and $\mathbf{X}^{R}_{i}$ is injected into the SE input to steer attention toward critical semantic regions, improving fusion accuracy. To capture multi-scale representations, the channel dimension is evenly split into four heads, each independently applying Low-Rank Adaptation.
@@
$\mathcal{L}=\underbrace{\|\hat{\mathbf{Z}}-\mathbf{Z}\|_{1}}_{\mathcal{L}_{1}}-\lambda_{\mathrm{inv}}\cdot\underbrace{\log\left|\det\left(\frac{\partial\bar{\mathbf{Z}}}{\partial\mathbf{X}}\right)\right|}_{\mathcal{L}_{inv}}+\lambda_{\mathrm{cos}}\cdot\underbrace{\left(1-\cos(\bar{\mathbf{Z}},\hat{\mathbf{Z}})\right)}_{\mathcal{L}_{cos}}$ (4)
@@
[…]
@@
We conduct experiments on three widely used hyperspectral datasets: Chikusei (Chikusei), PaviaU, and Houston. The Chikusei image ($2517\times 2335$, 128 bands) is split by using a $1000\times 2000$ top-left region for training and the remaining area, divided into $680\times 680$ patches, for testing. PaviaU ($610\times 340$, 103 bands) uses the top $340\times 340$ region for testing and the rest for training. For Houston ($349\times 1905$, 144 bands), the left $349\times 349$ area is used for testing. Following (simData), HRMSI are simulated, and LRHSI are obtained via Gaussian blur (kernel size $3\times 3$, std 0.5) and downsampling. Training samples are cropped into $64\times 64$ ground-truth patches paired with corresponding LRHSI ($16\times 16$) and HRMSI patches.
@@
For quantitative evaluation, we adopt four commonly-used metrics: Peak Signal-to-Noise Ratio (PSNR), Spectral Angle Mapper (SAM) (sam), Relative Global Dimensional Synthesis Error (ERGAS) (ERGAS), and Structural Similarity Index (SSIM) (SSIM).